# Feasibility of Learning Weighted Automata on a Semiring

### v4

@@ 1. Introduction
Our approach is somewhat different to what has been done so far: we look globally at all functions computed by finite state weighted automata and investigate, for each of them, whether they can potentially be learned by an Angluin-style algorithm, i.e. where the automata that are guessed (hypothesis automata, see Definition 2.10) are constructed in a prescribed way using finitely many membership queries (to be explained in detail below). Working over an arbitrary semiring, we define levels of guessability and a hierarchy of functions with respect to these levels: from the ones for which no such guess will be correct to the ones where there is a simple theoretical strategy to construct a correct guess. Our main result in this regard (Theorem 3.6) is a syntactic condition on the functions that have the potential to be learnt. For such a function (weakly guessable functions, see Definition 3.1) there must exist a finite state automaton of a prescribed type computing it: we call these automata literal (Defintion 3.2) and they are closely related to residual automata. In cases already widely studied in the literature (e.g. where the weights are drawn from fields or integral domains), it turns out that every function computed by a finite state automaton can be computed by a literal automaton. Thus, in the existing literature, focus is (reasonably) on algorithmic issues. In this paper, we focus on picturing the landscape for active learning in arbitrary semirings, and showing the limitations; for instance over the non-negative integers, it is easily seen that the example of van Heerdt, Kupke, Rot and Silva mentioned in the previous paragraph cannot be computed by a literal automaton.
@@ 2.2. Learning with membership and equivalence queries
 (2) That $X$ left-generates $\langle\varepsilon\rangle$ is completely clear from definition of row generating set. Since $X$ is a subset of the rows of the Hankel matrix of $f$, each element of $X$ may be written in the form $\langle w\rangle$ for some $w\in\Sigma^{*}$, in which case $wa\in\Sigma^{*}$ for all $a\in\Sigma$, and hence $\langle w\rangle\cdot a=\langle wa\rangle$ is left-generated by $X$. ∎
Example 2.5.
@@
We first set up some notation to describe the construction of hypothesis automaton.
Notation 2.6.
Given a function $f$ (and in particular the row indexed by $\varepsilon$ in its Hankel matrix), a finite non-empty subset $Q$ of vectors of $S^{\Sigma^{*}}$ and a non-empty subset $T$ of $\Sigma^{*}$, we define the left solution set $\Lambda_{Q,T}$ to be the set of vectors $\underline{\lambda}\in S^{Q\cup(Q\times\Sigma\times Q)}$ satisfying:
@@
Remark 2.7.
An element $\underline{\lambda}\in\Lambda_{Q,T}$ represents a collection of coefficients witnessing that $\langle\varepsilon\rangle_{T}$ and each of the $(\underline{q}\cdot a)_{T}$ (for all $\underline{q}\in Q$ and all $a\in\Sigma$) is a left-linear combinations of the $\underline{q}_{T}$ for $q$ in $Q$. In other words, $\Lambda_{Q,T}$ represents the solution space to the system of $S$-linear equations outlined in Notation 2.6. The key idea here is that the set $\Lambda_{Q,T}$ is non-empty if and only if a solution to this system exists; in the language introduced above, if and only if $Q$ is left-closed on $T$ and $\langle\varepsilon\rangle_{T}$ is left-generated by $Q$ restricted to $T$.
Remark 2.8.
@@
Proof.
Since each element of $Q$ is a row of the Hankel matrix, there exists a finite set of words $W\subseteq\Sigma^{*}$ such that $Q=\{\langle w\rangle:w\in W\}$. Since $\Lambda_{Q^{\prime}}\neq\emptyset$, we have that $Q^{\prime}$ is left-closed and left-generates the row $\langle\varepsilon\rangle$, and so $Q^{\prime}$ is row-generating for $f$, by Lemma 2.4. In particular, $\langle\varepsilon\rangle$ and each row $\langle wa\rangle$ where $w\in W$ and $a\in\Sigma$ can be expressed as a left linear combination of the elements of $Q^{\prime}$, and hence also of $Q$. Since $\langle wa\rangle=\langle w\rangle\cdot a$, it then follows easily that the system of equations given in Notation 2.6 has a solution. ∎
Definition 2.10.
@@
Fliess’s theorem [21] states that a function is computed by a weighted automaton over the field of reals if and only if its Hankel matrix has finite rank, and moreover, this rank is exactly the minimal number of states of a weighted automaton computing the function. There are many equivalent definitions of the rank of a matrix with entries over a field whose analogues cease to coincide over an arbitrary semirings, however, the proof of Fliess can be easily adapted to obtain the following statement, and since the details are instructive, we include a short proof.
Theorem 2.12.
@@
Before presenting the proof, we alert the reader to the fact that in Theorem 2.12 the set $Q$ is not assumed to be a set of rows of the Hankel matrix of $f$, and as we shall see below, in general, a function computed by a finite-state weighted automaton over an arbitrary semiring may or may not satisfy $\Lambda_{Q}\neq\emptyset$ for some finite subset $Q$ of the rows of its Hankel matrix. However, in view of Notation 2.6, Remark 2.7 and Definition 2.10, if there exists a finite set $Q$ of rows of the Hankel matrix such that $\Lambda_{Q}$ is non-empty, then for any finite set $T\subseteq\Sigma^{*}$ and any $\underline{\lambda}\in\Lambda_{Q}$ we have $\underline{\lambda}\in\Lambda_{Q}\subseteq\Lambda_{Q,T}$, so that (by construction) the hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ constructed from the finite $(Q,T)$ submatrix of the Hankel matrix of $f$, will be equal to the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ and, so by Theorem 2.12, will compute $f$. The success of Angluin-type algorithms in the existing literature is therefore typically based on the ability to find (using finitely many membership and equivalence queries): a finite set of words $Q$ such that $\Lambda_{Q}$ is non-empty, together with a finite set of words $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$ in the given setting.
Proof.
@@
Suppose that $Q\subseteq S^{\Sigma^{*}}$ is a finite subset such that $\Lambda_{Q}\neq\emptyset$ (or equivalently by Remark 2.7, $Q$ is left-closed and left-generates $\langle\varepsilon\rangle$) and let $\underline{\lambda}\in\Lambda_{Q}$. We shall show that $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$. To begin with, for each $\underline{q}$ in $Q$, let $\mathcal{H}_{\underline{q}}$ be the automaton obtained from $\mathcal{H}_{Q,\underline{\lambda}}$ by imposing the condition that $\underline{q}$ is the unique initial state with weight $1_{S}$, and keeping all other weights the same. We claim that for all $w$ in $\Sigma^{*}$, $\underline{q}_{w}$ is equal to the weight computed by $\mathcal{H}_{\underline{q}}$ on input $w$. The proof is by induction on the length of $w$. If $w=\varepsilon$, then $\underline{q}_{\varepsilon}$ is (by construction) equal to the final weight of $\underline{q}$ in $\mathcal{H}_{Q,\underline{\lambda}}$ which proves the property. For $a$ in $\Sigma$ and $w$ in $\Sigma^{*}$, $\underline{q}_{aw}=(\underline{q}\cdot a)_{w}=\bigoplus_{\underline{p}\in Q}\underline{\lambda}_{\underline{q},a,\underline{p}}\otimes\underline{p}_{w}$ since $\underline{\lambda}\in\Lambda_{Q}$. By induction hypothesis, $\underline{p}_{w}$ is the weight computed by $\mathcal{H}_{\underline{p}}$ on input $w$, and moreover $\underline{\lambda}_{\underline{q},a,\underline{p}}$ is (by construction) the weight of the transition from state $\underline{q}$ to state $\underline{p}$ labelled by $a$. So $\underline{q}_{aw}$ is indeed the weight computed by $\mathcal{H}_{\underline{q}}$ on input $aw$. To conclude the proof, recall that $\langle\varepsilon\rangle$ denotes the row of $\varepsilon$ in the Hankel matrix of $f$, and so by definition of the Hankel matrix of $f$ and the fact that $\underline{\lambda}\in\Lambda_{Q}$ we have $f(w)=\langle\varepsilon\rangle_{w}=\bigoplus_{\underline{q}\in Q}\underline{\lambda}_{q}\otimes\underline{q}_{w}$; by our previous observation (together with the fact that $\underline{\lambda}_{Q}$ is the initial state vector of $\mathcal{H}_{Q,\underline{\lambda}}$), this is then exactly the value computed by $\mathcal{H}_{Q,\lambda}$ on input $w$.
@@
where $\nu_{p}$ is the weight of the transition labelled by $a$ in $\mathcal{A}$ from $q$ to $p$. Let $w$ be a word in $\Sigma^{*}$ and as before let $\langle w\rangle$ denote the infinite row of the Hankel matrix of $f$ indexed by $w$. Let $\mu_{q}(w)$ be the value computed by $\mathcal{A}_{q}$ on input $w$. Then, by definition, $\langle w\rangle=\bigoplus_{q\in Q}\mu_{q}(w)\otimes\underline{q}$, showing that $\underline{Q}$ is row-generating for $f$ and hence in particular $\langle\varepsilon\rangle$ is left-generated by $\underline{Q}$. ∎
Remark 2.14.
@@
The previous result together with its dual (see Theorem A.3 of the appendix) give the following:
Corollary 2.15.
@@
In Section 3, we define the classes of weakly guessable, guessable and strongly guessable functions and give several equivalent characterisations for them in Theorem 3.6 and 3.13. In Section 4, we give a general hierarchy for these classes with some specific examples, and give conditions for this hierarchy to collapse. This allows us to have a clearer picture of what happens for several well-studied examples of semirings arising in application areas, where limitations of this approach may be of interest. This opens many questions for further research that are stated in the concluding Section 5.
@@ 3.1. Weakly guessable functions
Observe that by Theorem 2.12, this implies that for all $\underline{\lambda}\in\Lambda_{Q}$ the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and hence for any non-empty set of words $T$, there exists an hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ computing $f$ since (by Remark 2.8) $\Lambda_{Q}\subseteq\Lambda_{Q,T}$. So this definition means that there is “some hope” that an algorithm à la Angluin can find a correct automaton, but does not really give a handle on what these functions look like. The following definition (literal automaton) turns out to provide a syntactic charaterisation of these functions in terms of automata, and links to residual automata. A literal automaton is a residual automaton, as defined for example for non-deterministic or nominal automata [22, 27] with extra conditions to deal with the weights. It is very close to the reduced residual automata in the stochastic settings defined in [17], with a slight difference in the way the weights are handled - the reason being that, while non-zero probabilities have an inverse for the multiplication, this is not systematically the case in general semirings. We use the terminology literal automata to avoid any confusion with [17], our choice of terminology reflecting the fact that each state of a literal automaton can be labelled by a word with the property that any run from the initial state labelled by that word terminates in the given state.
Definition 3.2.
@@
Remark 3.3.
In a literal automaton $\mathcal{A}$ with set of states $Q$, each state $q$ can be labelled by a word $\sigma(q)$ which when taken as input terminates in state $q$ (via a unique run, which has weight $1_{S}$), and so in this case the set $\underline{Q}=\{\underline{q}\mid q\text{ is a state of }\mathcal{A}\}$ defined in Theorem 2.12 is easily seen to be precisely the set of rows of the Hankel matrix of $\mathcal{A}$ indexed by the words chosen to label the states, that is, $\underline{Q}=\{\langle\sigma(q)\rangle:q\in Q\}$.
@@
Remark 3.7.
Before presenting the proof, it is instructive to compare Theorem 3.6 (characterising weakly guessable functions) with Theorem 2.12 (characterising functions computed by a finite weighted automata). Notice that point (3) of Theorem 3.6 states that there is a finite subset of the rows of the Hankel matrix that left-generates every row of the Hankel matrix, whilst Theorem 2.12, states that the rows of the Hankel matrix can be generated by finitely many elements of $S^{\Sigma^{*}}$ (which need not themselves be rows of the Hankel matrix; see also Remark 2.13).
Proof.
(1. implies 2.) Since $f$ is weakly guessable there is a finite set of rows $Q$ of the Hankel matrix such that $\Lambda_{Q}$ is non-empty. Let $W$ be a finite set of words with $|W|=|Q|$ such that $Q$ is the set of rows of the Hankel matrix indexed by $W$ and let $W^{\prime}$ denote the set of all prefixes of words in $W$, with $Q^{\prime}$ the corresponding set of rows of the Hankel matrix, so that $Q\subseteq Q^{\prime}=\{\langle q\rangle\mid q\in W^{\prime}\}$. Since $\Lambda_{Q}\neq\emptyset$, we have seen (in Lemma 2.9) that $\Lambda_{Q^{\prime}}\neq\emptyset$; we aim to show that there exists $\underline{\mu}\in\Lambda_{Q^{\prime}}$ such that the corresponding Hankel automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ (which by Theorem 2.12 computes $f$) is literal. To this end, we let $\underline{\lambda}\in\Lambda_{Q}$ so that (by Theorem  2.12) $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and use the values in $\underline{\lambda}$ to define $\underline{\mu}\in S^{Q^{\prime}\cup(Q^{\prime}\times\Sigma\times Q^{\prime})}$ as follows:
@@
By construction we then have that $\underline{\mu}$ belongs to $\Lambda_{Q^{\prime}}$ and hence by Lemma 2.12, $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ computes $f$. The automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ is also literal by construction, which concludes the proof.
(2. implies 3.) Let $\mathcal{A}$ be a literal automaton computing $f$ and $Q$ its set of states. By Theorem 2.12, the set $\underline{Q}=\{\underline{q}\mid q\in Q\}$, where $\underline{q}_{w}$ is defined as the value computed by the automaton $\mathcal{A}_{q}$ on input $w$, is row-generating for $f$ and, as observed in Remark 3.3, there is a finite set of words $\{\sigma(q)\mid q\in Q\}$ such that $\underline{Q}=\{\langle\sigma(q)\rangle\mid q\in Q\}$.
@@
(4. implies 1.) For all positive integers $i$, let $X_{i}$ be the left-semimodule generated by the set $Q_{i}$ containing all rows of $F$ indexed by words of length at most $i$. Then the $X_{i}$ satisfy the hypothesis of the weak acending chain condition and hence there is $n$ such that for all $m$, $X_{m}=X_{n}$. Necessarily the set $Q_{n}$ is row-generating and hence (by Lemma 2.4) left-closed, and left-generates $\langle\varepsilon\rangle$, so there exists some $\lambda\in\Lambda_{Q}$, giving $\Lambda_{Q_{n}}\neq\emptyset$ wich demonstrates that $f$ is weakly guessable and concludes the proof. ∎
Remark 3.8.
As noted in Remark 2.14, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions which we refer to as weakly co-guessable, column generating sets, and weak co-ascending chain condition: see Appendix A for full details. We say that an automaton is co-literal if its mirror is literal. With these definitions in place, one can prove the following entirely analogous result.
Theorem 3.9.
@@
Remark 3.10.
The semirings where all functions computed by finite state weighted automata satisfy the weak ascending and co-ascending chain conditions are therefore the semirings for which all functions computed by finite state weighted automata are both weakly guessable and weakly co-guessable (by Theorem 3.6 and its dual Theorem 3.9). This means that over such a semiring all functions computed by a finite weighted automaton have the potential to be learnt, provided one can make a lucky guess for $Q$, $T$ and $\lambda$. This is in particular the case for $\mathbb{B}$ (or indeed any finite semiring), $\mathbb{R}$ (or indeed, any field), and $\mathbb{Z}$ (or indeed, any principal ideal domain c.f.  [36]). More generally, if $R$ is any left-Noetherian ring, then all finitely generated left $R$-modules are Noetherian, and hence by Remark 3.5, all functions computed by finite state weighted automata over $R$ are weakly guessable. This covers the case of finite rings, fields, principal ideal domains, and division rings.

### v5

@@ 1. Introduction
Our approach is somewhat different to what has been done so far: we look globally at all functions computed by finite state weighted automata and investigate, for each of them, whether they can potentially be learned by an Angluin-style algorithm, i.e. where the automata that are guessed (hypothesis automata, see Definition 2.2) are constructed in a prescribed way using finitely many membership queries (to be explained in detail below). Working over an arbitrary semiring, we define levels of guessability and a hierarchy of functions with respect to these levels: from the ones for which no such guess will be correct to the ones where there is a simple theoretical strategy to construct a correct guess. Our main result in this regard (Theorem 12) is a syntactic condition on the functions that have the potential to be learnt. For such a function (weakly guessable functions, see Definition 3.1) there must exist a finite state automaton of a prescribed type computing it: we call these automata literal (Definition 3.1) and they are closely related to residual automata. In cases already widely studied in the literature (e.g. where the weights are drawn from fields or integral domains), it turns out that every function computed by a finite state automaton can be computed by a literal automaton. Thus, in the existing literature, focus is (reasonably) on algorithmic issues. In this paper, we focus on picturing the landscape for active learning in arbitrary semirings, and showing the limitations; for instance over the non-negative integers, it is easily seen that the example of van Heerdt, Kupke, Rot and Silva mentioned in the previous paragraph cannot be computed by a literal automaton.
@@ 2.2. Learning with membership and equivalence queries
 (2) That $X$ left-generates $\langle\varepsilon\rangle$ is completely clear from definition of row generating set. Since $X$ is a subset of the rows of the Hankel matrix of $f$, each element of $X$ may be written in the form $\langle w\rangle$ for some $w\in\Sigma^{*}$, in which case $wa\in\Sigma^{*}$ for all $a\in\Sigma$, and hence $\langle w\rangle\cdot a=\langle wa\rangle$ is left-generated by $X$.
{exa}
@@
We first set up some notation to describe the construction of hypothesis automaton. {nota} Given a function $f$ (and in particular the row indexed by $\varepsilon$ in its Hankel matrix), a finite non-empty subset $Q$ of vectors of $S^{\Sigma^{*}}$ and a non-empty subset $T$ of $\Sigma^{*}$, we define the left solution set $\Lambda_{Q,T}$ to be the set of vectors $\underline{\lambda}\in S^{Q\cup(Q\times\Sigma\times Q)}$ satisfying:
@@
Remark 2.
An element $\underline{\lambda}\in\Lambda_{Q,T}$ represents a collection of coefficients witnessing that $\langle\varepsilon\rangle_{T}$ and each of the $(\underline{q}\cdot a)_{T}$ (for all $\underline{q}\in Q$ and all $a\in\Sigma$) is a left-linear combinations of the $\underline{q}_{T}$ for $q$ in $Q$. In other words, $\Lambda_{Q,T}$ represents the solution space to the system of $S$-linear equations outlined in Notation 2.2. The key idea here is that the set $\Lambda_{Q,T}$ is non-empty if and only if a solution to this system exists; in the language introduced above, if and only if $Q$ is left-closed on $T$ and $\langle\varepsilon\rangle_{T}$ is left-generated by $Q$ restricted to $T$.
Remark 3.
@@
Proof 2.2.
Since each element of $Q$ is a row of the Hankel matrix, there exists a finite set of words $W\subseteq\Sigma^{*}$ such that $Q=\{\langle w\rangle:w\in W\}$. Since $\Lambda_{Q^{\prime}}\neq\emptyset$, we have that $Q^{\prime}$ is left-closed and left-generates the row $\langle\varepsilon\rangle$, and so $Q^{\prime}$ is row-generating for $f$, by Lemma 1. In particular, $\langle\varepsilon\rangle$ and each row $\langle wa\rangle$ where $w\in W$ and $a\in\Sigma$ can be expressed as a left linear combination of the elements of $Q^{\prime}$, and hence also of $Q$. Since $\langle wa\rangle=\langle w\rangle\cdot a$, it then follows easily that the system of equations given in Notation 2.2 has a solution.
{defi}
@@
Fliess’s theorem [21] states that a function is computed by a weighted automaton over the field of reals if and only if its Hankel matrix has finite rank, and moreover, this rank is exactly the minimal number of states of a weighted automaton computing the function. There are many equivalent definitions of the rank of a matrix with entries over a field whose analogues cease to coincide over an arbitrary semirings, however, the proof of Fliess’ Theorem can be easily adapted to obtain the following statement, and since the details are instructive, we include a short proof.
Theorem 5.
@@
Before presenting the proof, we alert the reader to the fact that in Theorem 5 the set $Q$ is not assumed to be a set of rows of the Hankel matrix of $f$, and as we shall see below, in general, a function computed by a finite-state weighted automaton over an arbitrary semiring may or may not satisfy $\Lambda_{Q}\neq\emptyset$ for some finite subset $Q$ of the rows of its Hankel matrix. However, in view of Notation 2.2, Remark 2 and Definition 2.2, if there exists a finite set $Q$ of rows of the Hankel matrix such that $\Lambda_{Q}$ is non-empty, then for any finite set $T\subseteq\Sigma^{*}$ and any $\underline{\lambda}\in\Lambda_{Q}$ we have $\underline{\lambda}\in\Lambda_{Q}\subseteq\Lambda_{Q,T}$, so that (by construction) the hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ constructed from the finite $(Q,T)$ submatrix of the Hankel matrix of $f$, will be equal to the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ and, so by Theorem 5, will compute $f$. The success of Angluin-type algorithms in the existing literature is therefore typically based on the ability to find (using finitely many membership and equivalence queries): a finite set of words $Q$ such that $\Lambda_{Q}$ is non-empty, together with a finite set of words $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$ in the given setting.
Proof 2.3.
@@
Suppose that $Q\subseteq S^{\Sigma^{*}}$ is a finite subset such that $\Lambda_{Q}\neq\emptyset$ (or equivalently by Remark 2, $Q$ is left-closed and left-generates $\langle\varepsilon\rangle$) and let $\underline{\lambda}\in\Lambda_{Q}$. We shall show that $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$. To begin with, for each $\underline{q}$ in $Q$, let $\mathcal{H}_{\underline{q}}$ be the automaton obtained from $\mathcal{H}_{Q,\underline{\lambda}}$ by imposing the condition that $\underline{q}$ is the unique initial state with weight $1_{S}$, and keeping all other weights the same. We claim that for all $w$ in $\Sigma^{*}$, $\underline{q}_{w}$ is equal to the weight computed by $\mathcal{H}_{\underline{q}}$ on input $w$. The proof is by induction on the length of $w$. If $w=\varepsilon$, then $\underline{q}_{\varepsilon}$ is (by construction) equal to the final weight of $\underline{q}$ in $\mathcal{H}_{Q,\underline{\lambda}}$ which proves the property. For $a$ in $\Sigma$ and $w$ in $\Sigma^{*}$, $\underline{q}_{aw}=(\underline{q}\cdot a)_{w}=\bigoplus_{\underline{p}\in Q}\underline{\lambda}_{\underline{q},a,\underline{p}}\otimes\underline{p}_{w}$ since $\underline{\lambda}\in\Lambda_{Q}$. By induction hypothesis, $\underline{p}_{w}$ is the weight computed by $\mathcal{H}_{\underline{p}}$ on input $w$, and moreover $\underline{\lambda}_{\underline{q},a,\underline{p}}$ is (by construction) the weight of the transition from state $\underline{q}$ to state $\underline{p}$ labelled by $a$. So $\underline{q}_{aw}$ is indeed the weight computed by $\mathcal{H}_{\underline{q}}$ on input $aw$. To conclude the proof, recall that $\langle\varepsilon\rangle$ denotes the row of $\varepsilon$ in the Hankel matrix of $f$, and so by definition of the Hankel matrix of $f$ and the fact that $\underline{\lambda}\in\Lambda_{Q}$ we have $f(w)=\langle\varepsilon\rangle_{w}=\bigoplus_{\underline{q}\in Q}\underline{\lambda}_{q}\otimes\underline{q}_{w}$; by our previous observation (together with the fact that $\underline{\lambda}_{Q}$ is the initial state vector of $\mathcal{H}_{Q,\underline{\lambda}}$), this is then exactly the value computed by $\mathcal{H}_{Q,\lambda}$ on input $w$.
@@
where $\nu_{p}$ is the weight of the transition labelled by $a$ in $\mathcal{A}$ from $q$ to $p$. Let $w$ be a word in $\Sigma^{*}$ and as before let $\langle w\rangle$ denote the infinite row of the Hankel matrix of $f$ indexed by $w$. Let $\mu_{q}(w)$ be the value computed by $\mathcal{A}_{q}$ on input $w$. Then, by definition, $\langle w\rangle=\bigoplus_{q\in Q}\mu_{q}(w)\otimes\underline{q}$, showing that $\underline{Q}$ is row-generating for $f$ and hence in particular $\langle\varepsilon\rangle$ is left-generated by $\underline{Q}$.
Remark 7.
@@
The previous result together with its dual (see Theorem 36 of the appendix) give the following:
Corollary 8.
@@
In Section 3, we define the classes of weakly guessable, guessable and strongly guessable functions and give several equivalent characterisations for them in Theorem 12 and 17. In Section 4, we give a general hierarchy for these classes with some specific examples, and give conditions for this hierarchy to collapse. This allows us to have a clearer picture of what happens for several well-studied examples of semirings arising in application areas, where limitations of this approach may be of interest. This opens many questions for further research that are stated in the concluding Section 5.
@@ 3.1. Weakly guessable functions
Observe that by Theorem 5, this implies that for all $\underline{\lambda}\in\Lambda_{Q}$ the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and hence for any non-empty set of words $T$, there exists an hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ computing $f$ since (by Remark 3) $\Lambda_{Q}\subseteq\Lambda_{Q,T}$. So this definition means that there is “some hope” that an algorithm à la Angluin can find a correct automaton, but does not really give a handle on what these functions look like. The following definition (literal automaton) turns out to provide a syntactic characterisation of these functions in terms of automata, and links to residual automata. A literal automaton is a residual automaton, as defined for example for non-deterministic or nominal automata [22, 27] with extra conditions to deal with the weights. It is very close to the reduced residual automata in the stochastic settings defined in [17], with a slight difference in the way the weights are handled - the reason being that, while non-zero probabilities have an inverse for the multiplication, this is not systematically the case in general semirings. We use the terminology literal automata to avoid any confusion with [17], our choice of terminology reflecting the fact that each state of a literal automaton can be labelled by a word with the property that any run from the initial state labelled by that word terminates in the given state.
{defi}
@@
Remark 10.
In a literal automaton $\mathcal{A}$ with set of states $Q$, each state $q$ can be labelled by a word $\sigma(q)$ which when taken as input terminates in state $q$ (via a unique run, which has weight $1_{S}$), and so in this case the set $\underline{Q}=\{\underline{q}\mid q\text{ is a state of }\mathcal{A}\}$ defined in Theorem 5 is easily seen to be precisely the set of rows of the Hankel matrix of $\mathcal{A}$ indexed by the words chosen to label the states, that is, $\underline{Q}=\{\langle\sigma(q)\rangle:q\in Q\}$.
@@
Remark 13.
Before presenting the proof, it is instructive to compare Theorem 12 (characterising weakly guessable functions) with Theorem 5 (characterising functions computed by a finite weighted automata). Notice that point (3) of Theorem 12 states that there is a finite subset of the rows of the Hankel matrix that left-generates every row of the Hankel matrix, whilst Theorem 5, states that the rows of the Hankel matrix can be generated by finitely many elements of $S^{\Sigma^{*}}$ (which need not themselves be rows of the Hankel matrix; see also Remark 6).
Proof 3.1.
(1. implies 2.) Since $f$ is weakly guessable there is a finite set of rows $Q$ of the Hankel matrix such that $\Lambda_{Q}$ is non-empty. Let $W$ be a finite set of words with $|W|=|Q|$ such that $Q$ is the set of rows of the Hankel matrix indexed by $W$ and let $W^{\prime}$ denote the set of all prefixes of words in $W$, with $Q^{\prime}$ the corresponding set of rows of the Hankel matrix, so that $Q\subseteq Q^{\prime}=\{\langle q\rangle\mid q\in W^{\prime}\}$. Since $\Lambda_{Q}\neq\emptyset$, we have seen (in Lemma 4) that $\Lambda_{Q^{\prime}}\neq\emptyset$; we aim to show that there exists $\underline{\mu}\in\Lambda_{Q^{\prime}}$ such that the corresponding Hankel automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ (which by Theorem 5 computes $f$) is literal. To this end, we let $\underline{\lambda}\in\Lambda_{Q}$ so that (by Theorem  5) $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and use the values in $\underline{\lambda}$ to define $\underline{\mu}\in S^{Q^{\prime}\cup(Q^{\prime}\times\Sigma\times Q^{\prime})}$ as follows:
@@
By construction we then have that $\underline{\mu}$ belongs to $\Lambda_{Q^{\prime}}$ and hence by Lemma 5, $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ computes $f$. The automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ is also literal by construction, which concludes the proof.
(2. implies 3.) Let $\mathcal{A}$ be a literal automaton computing $f$ and $Q$ its set of states. By Theorem 5, the set $\underline{Q}=\{\underline{q}\mid q\in Q\}$, where $\underline{q}_{w}$ is defined as the value computed by the automaton $\mathcal{A}_{q}$ on input $w$, is row-generating for $f$ and, as observed in Remark 10, there is a finite set of words $\{\sigma(q)\mid q\in Q\}$ such that $\underline{Q}=\{\langle\sigma(q)\rangle\mid q\in Q\}$.
@@
(4. implies 1.) For all positive integers $i$, let $X_{i}$ be the left-semimodule generated by the set $Q_{i}$ containing all rows of $F$ indexed by words of length at most $i$. Then the $X_{i}$ satisfy the hypothesis of the weak ascending chain condition and hence there is $n$ such that for all $m$, $X_{m}=X_{n}$. Necessarily the set $Q_{n}$ is row-generating and hence (by Lemma 1) left-closed, and left-generates $\langle\varepsilon\rangle$, so there exists some $\lambda\in\Lambda_{Q}$, giving $\Lambda_{Q_{n}}\neq\emptyset$ which demonstrates that $f$ is weakly guessable and concludes the proof.
Remark 14.
As noted in Remark 7, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions which we refer to as weakly co-guessable, column generating sets, and weak co-ascending chain condition: see Appendix A for full details. We say that an automaton is co-literal if its mirror is literal. With these definitions in place, one can prove the following entirely analogous result.
Theorem 15.
@@
Remark 16.
The semirings where all functions computed by finite state weighted automata satisfy the weak ascending and co-ascending chain conditions are therefore the semirings for which all functions computed by finite state weighted automata are both weakly guessable and weakly co-guessable (by Theorem 12 and its dual Theorem 15). This means that over such a semiring all functions computed by a finite weighted automaton have the potential to be learnt, provided one can make a lucky guess for $Q$, $T$ and $\lambda$. This is in particular the case for $\mathbb{B}$ (or indeed any finite semiring), $\mathbb{R}$ (or indeed, any field), and $\mathbb{Z}$ (or indeed, any principal ideal domain c.f.  [36]). More generally, if $R$ is any left-Noetherian ring, then all finitely generated left $R$-modules are Noetherian, and hence by Remark 11, all functions computed by finite state weighted automata over $R$ are weakly guessable. This covers the case of finite rings, fields, principal ideal domains, and division rings.